Solve by finding an integrating factor

(x^2+y^2+2x)dx + 2ydy =0


1
Expert's answer
2021-04-15T07:20:54-0400

Solution


given the differential equation:

"(x^2+y^2+2x)dx+2ydy=0" ...(1)

let "f(x,y)=x^2+y^2+2x"

"\implies g(y)=2y"


so

f'"_y(x,y)=2y"

"g'(x)=0"


using Integration factor method,

"\frac{du}{dx}=\frac{(f_y-g_x)u}{g}"

"\frac{du}{dx}=\frac{(2y-0)u}{2y}"

"\frac{du}{dx}=\frac{(2y)u}{2y}"

"\frac{du}{dx}=u"

on solving we get integrated factor


"u=e^x"

Multiply equation 1 by integrated factor

"e^x(x^2+y^2+2x)dx+2ye^xdy=0"

let M="(x^2+y^2+2x)e^x"

N="2ye^x"

"\frac{dM}{dy}=2ye^x"

"\frac{dN}{dx}=2ye^x"


Hence "\frac{dM}{dy}=\frac{dN}{dx}"


Solution of above equation is given by,


Let I(x,y) be a implict function

I(x,y)="\int M(x,y)dx"

="\int e^x(x^2+y^2+2x)dx"


="x^2e^x-2xe^x+2e^x+y^2e^x+2xe^x-2e^x"

="x^2e^x+y^2e^x" +f(y)


Let differentiate I(x,y) with respect to y

"\frac{dI}{dy}=" N(x,y)

"2ye^x+\frac{df}{dy}=2ye^x"

"\frac{df}{dy}=0
=constant"

I(x,y)="x^2e^x+y^2e^x+c"


at "x=y=1"

"0=e+e+c\\

c+2e=0\\

c=-2e"


hence this is the required solution is;

"x^2e^x+y^2e^x-2e=0"

"x^2+y^2=2e^{1-x}" .